hree identical metal plates with large surface areas are kept parallel to each other.The leftmost plate is given a charge Q, the rightmost is given a charge -2Q & the middle one remains neutral.Find the charge appearing on the outer surface of the rightmost plate.

Grade:

1 Answers

madhav purohit
11 years ago

Consider the gaussian surface the induced charge be as show in the fig. [in the quest.]the net field at p due to all the charge is zero 


-2Q+q/2Ae0 (left)+q/2Ae0

+q-Q+q=0

q=3/2Q

charge on the right side of right most plate

=-2Q+q=-2Q+3/2Q

=-Q/2

and this is the solution and if u hve any other doubt about physics and contact me http://www.facebook.com/madhav.purohit.984